# Hyperparameter Optimization in Machine Learning

### v1

@@ 8.4.2 Meta-learning as bilevel programming
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO is a fast-moving field, and the current landscape is changing frequently, so the provided list here is by no means exhaustive and only represents a snapshot.
@@ 9.1 Open Source Libraries and Frameworks
One of the first libraries that provides Bayesian optimization for hyperparameter optimization was the Python package Spearmint [snoek_practical_2012], which implements Bayesian optimization using Gaussian processes (see Section 4.1). At the same time, bergstra2011algorithms provided HyperOpt, offering Bayesian hyperparameter optimization by using the tree-parzen estimator instead of Gaussian processes. Hutter2011:Sequential introduced SMAC, which implemented Bayesian optimization strategies using Random Forests. Initially targeting for more generic algorithm configuration problems, such as configuring SAT solvers, it also offered functionality for HPO of machine learning methods. While the initial version of SMAC was written in Java, there is also a newer version fully implemented in Python [smac3], that also provides more advanced functionalities, such as multi-fidelity optimization (see Chapter 5).
Some of the first Python libraries that implemented multi-fidelity HPO, such as Fabolas [klein2017fast] or Hyperband [li2016hyperband], was the RoBO library by klein-bayesopt17 and Dragonfly by kandasamy-jmlr19.
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on the large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distributed the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
Recent pytorch implementations of gradient-based hyperparameter optimziation algorithms include the libraries HyperTorch [grazzi2020iteration], higher [grefenstette2019generalized] and Betty [choe2023betty]. JaxOpt [blondel2022efficient] is a Jax implementation that focuses on implicit differentiation.
@@ 9.3 Benchmarking
A key challenge for an empirical driven research field such as HPO to make progress is the thorough evaluation of new approaches and comparison to baseline methods. However, due to the expensive nature of the underlying optimization problem (see Chapter 2), it is often prohibitively expensive to obtain statistically reliable results including multiple repetitions, which are necessary to characterize HPO algorithms that often come with intrinsic randomness.
@@
Hyperparameter optimization is a mature field with well-established algorithmic solutions and industry-level software frameworks available. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (model-based, multi-fidelity, population-based, and gradient-based) and briefly highlighted connections to closely related areas such neural architecture search and meta-learning. There remain open issues and interesting directions for improvement which we have not covered and briefly sketch below.
Response functions for unsupervised learning.
@@
HPO requires a well defined response function to be minimized (see Section 2.2), which in turn requires a well defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning. A notable example is anomaly detection, an area rich of methods, results, and applications [pang:2021:deep_learning_anomaly]. The anomaly or novelty detection problem essentially consists of identifying the support of the distribution behind the data, so that low-probability points can be detected as outliers. For this problem, [ma:2023:need_unsupervised_outlier] survey a number of internal (i.e., not using any ground truth labels) approaches for assessing outlier detection models and conclude that “none of the existing and adapted strategies would be practically useful.”
@@
A second important example is generative models. In this case the goal is to construct a model of the data distribution such that sampling from that model produces fake data points that are indistinguishable from real data. Again, the literature is rich in terms of models, ranging from early generative adversarial networks [DBLP:journals/corr/GoodfellowPMXWOCB14] to normalizing flows [pmlr-v37-rezende15] and more recently probabilistic diffusion models [ho:2020:denoising_diffusion_probabilistic]. In the particular case of images, popular metrics used to compare different generative models are the Inception Score [NIPS2016_8a3363ab] and the Fréchet inception distance (FID) [DBLP:conf/nips/HeuselRUNH17]. These metrics only capture one aspect of the generative process, namely fidelity. However, the problem of assessing the quality of a model is inherently more complicated and multidimensional: a perfectly faithful model that only produces samples that are too close to training data points is useless (implying an additional dimension related to diversity or creativity). There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports (of real and fake data) are estimated and compared (fake points off the support of real points are interpreted as false positives, and real points out of the support of fake points are interpreted as false negatives). The idea has sparkled a number of practically computable metrics that [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to optimize hyperparameters. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions would require reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionalty issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced in [yang:2020:evaluation_generative_models] for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data.
@@
Learning to (hyperparameter) optimize
In section 8.4 we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning an hyperparamter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and a response function of a (or several) standard learning algorithm. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g. validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on experience optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves, which we discussed in Chapter 5 when covering multi-fidelity methods; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modelling and learning to optimize, a sub-field of meta-learning, [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
The usage of rich meta-data, past experience, and leveraging similarities across standard learning algorithms may hold the promise of discovering HPO algorithm that are more efficient than classic approaches. However, significant challenges remain in developing truly general and robust meta-learning approaches for HPO, including handling the diversity of problem types, scalability to high-dimensional spaces, and transferability across vastly different domains and algorithms.
HPO and foundation models.
@@
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become standard paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices, which are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparamter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (cf. Chapter 7) that optimizes for discrete “trigger” tokens, Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; cf. Section 5.2.2); Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches (with discrete gradient estimators) may prove useful in these directions,
In addition, the emergence of agents and agent networks where various LLMs interact also by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs (APIs). These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the the communication protocols and the interactions graph between agents. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.

### v2

@@ 8.4.2 Meta-learning of common representations
8.5 Transfer of hyperparameters across model scale
A foundation model [brown2020language], such as a large language model, is a large-scale, pre-trained machine learning model, typically based on deep learning architectures like transformers [vaswani_attention_2017], that serves as a general-purpose backbone for various downstream tasks through fine-tuning or prompting. These models mark a cornerstone in artificial intelligence and have led to several breakthroughs in code generation, natural language processing, and common-sense reasoning. However, the expensive pre-training phase requires vast amounts of data and tremendous GPU resources.
While this also depends on a careful selection of hyperparameters, the hyperparameter optimization methods presented here are not directly applicable. Even efficient state-of-the-art multi-fidelity approaches (see Chapter 5) require at least a few training runs on the full budget, but for contemporary LLMs, this is infeasible.
Inspired by a theoretical perspective from tensor programs, a recent line of work [yang-arxiv22, everett-icml24a] approaches this problem from a different angle. While hyperparameters of neural networks, such as the learning rate, change when we increase the width or depth of the network and need to be adjusted accordingly, the idea here is to change the parameterization of neural networks, such that the optimal hyperparameters stay the same when scaling from smaller models to larger models. This allows for finding the optimal hyperparameters on a small scale and directly transferring them to the largest scale.
The original parameterization proposed by [yang-arxiv22], called $\mu$-parameterization, initially provided only a theoretical foundation for scaling model width, however they also provided empirical results confirming its effectiveness for depth scaling. Recent work by [everett-icml24a] has shown that even hyperparameters such as the learning rate can be transferred across different model widths using the standard parameterization of neural networks by adapting them on a per-layer basis.
While this line of research has shown promising results, it has so far been mostly theoretical and primarily focused on width scaling. Nevertheless, it has the potential to lead to more practical hyperparameter optimization methods that scale efficiently to large foundation models.
@@ 8.5 Transfer of hyperparameters across model scale
In this chapter, we provide an overview of different open-source libraries and commercial systems that implement methods discussed in the previous chapters. HPO software landscape changes frequently as new problems, methods and application domain become relevant: the provided list here is by no means exhaustive and only represents a snapshot.
@@ 9.1 Open Source Libraries and Frameworks
The Python package Spearmint [snoek_practical_2012] is one of the first libraries that provides Bayesian optimization for hyperparameter optimization. It implements Bayesian optimization using Gaussian processes (see Section 4.2). At the same time, bergstra2011algorithms provided HyperOpt, offering Bayesian hyperparameter optimization by using the tree Parzen estimator (Section 4.2.5) instead of Gaussian processes (Section 4.2.2). Hutter2011:Sequential introduced SMAC, which implemented Bayesian optimization strategies using Random Forests. Initially targeting at more generic algorithm configuration problems, such as configuring SAT solvers, it also offered functionality for HPO of machine learning methods. The current version [smac3] is written in Python and also provides more advanced functionalities, such as multi-fidelity optimization (see Chapter 5).
Among the first Python libraries that implemented multi-fidelity HPO, such as Fabolas [klein2017fast] or Hyperband [li2016hyperband], was the RoBO library by klein-bayesopt17 and Dragonfly by kandasamy-jmlr19.
While previous libraries, such as Spearmint, allowed to parallelize trials across processes on the same machine, more recent libraries focus more on the large-scale distributed setting to distribute the evaluation of trials across multiple instances or compute nodes. For example, RayTune [liaw2018tune], distributes the evaluation of hyperparameter configurations using the compute framework Ray. Another example is Orion, which supports both high-performance computing infrastructures and containerized environments. salinas-automl22 recently presented Syne Tune, which supports different backends that allow to distribute the HPO process either locally or on the cloud. Furthermore, it also provides a backend to simulate the evaluation of expensive objective functions using surrogate models (see Section 9.3). This allows to simulate large-scale asynchronous HPO on a single compute core.
Recent PyTorch implementations of gradient-based hyperparameter optimization algorithms include the libraries HyperTorch [grazzi2020iteration], higher [grefenstette2019generalized] and Betty [choe2023betty]. JaxOpt [blondel2022efficient] is a Jax implementation that focuses on implicit differentiation.
@@ 9.3 Benchmarking
A key challenge for an empirical driven research field such as HPO to make progress is the thorough evaluation of new approaches and comparison to baseline methods. However, due to the expensive nature of the underlying optimization problem, it is often prohibitively expensive to obtain statistically reliable results including multiple repetitions, which are necessary to characterize HPO algorithms that often come with intrinsic randomness.
@@
From some perspective, hyperparameter optimization may be considered by now a mature field with well-established algorithmic solutions and with industry-level software frameworks. It is a crucial ingredient for automated machine learning, an active research area that has even started a conference in 2022. In this monograph, we have reviewed in depth the major classes of methods (random and quasi-random search, model-based, multi-fidelity, population-based, and gradient-based) and highlighted connections to closely related areas such neural architecture search and meta-learning.
Yet, in all classes of algorithms there remain open issues and interesting directions for improvement. In addition, the development of hybrid algorithms has so far been under-explored, but could potentially bring notable benefits in practical applications. We do not discuss these particular aspects here, but rather touch upon some general topics which we feel may become (or rather, are already becoming) broad and important research directions in HPO.
10.1 Response functions for unsupervised learning
HPO requires a well defined response function to be minimized (see Section 2.2), which in turn requires a well defined performance metric. While suitable metrics are abundant for supervised learning, the same is not true for unsupervised learning.
@@ 10.1 Response functions for unsupervised learning
A notable example is anomaly detection, an area rich of methods, results, and applications [pang:2021:deep_learning_anomaly]. The anomaly or novelty detection problem essentially consists of identifying the high probability region of the distribution behind the data, so that points outside this region can be detected as outliers. For this problem, [ma:2023:need_unsupervised_outlier] survey a number of internal (i.e., not using any ground truth labels) approaches for assessing outlier detection models and conclude that “none of the existing and adapted strategies would be practically useful.”
@@
A second important example is generative models. In this case the goal is to construct a model of the data distribution such that sampling from that model produces fake data points that are indistinguishable from real data. Again, the literature is rich in terms of models, ranging from early generative adversarial networks [DBLP:journals/corr/GoodfellowPMXWOCB14] to normalizing flows [pmlr-v37-rezende15] and more recently probabilistic diffusion models [ho:2020:denoising_diffusion_probabilistic]. In the particular case of images, popular metrics used to compare different generative models are the Inception Score [NIPS2016_8a3363ab] and the Fréchet inception distance (FID) [DBLP:conf/nips/HeuselRUNH17]. These metrics only capture one aspect of the generative process, namely fidelity. However, the problem of assessing the quality of a model is inherently more complicated: it is not sufficient to generate faithful points, we also need to cover all regions where the data distribution is high.
There have been interesting attempts to interpret this perspective in terms of precision and recall [sajjadi:2018:assessing_generative_models, simon:2019:revisiting_precision_recall], where the two supports (of real and fake data) are estimated and compared (fake points off the support of real distribution are interpreted as false positives, and real points outside the support of the model’s distribution are interpreted as false negatives). The idea has sparkled a number of practically computable metrics [kynkaanniemi:2019:improved_precision_recall, naeem:2020:reliable_fidelity_diversitya, park:2023:probabilistic_precision_recall] that however have not been shown to be effective to guide HPO. A unidimensional metric that aims to assess the similarity between the model and the real data distributions is the overlapping area. Computing the overlaps between these two distributions requires reliable density estimation, which is however difficult for high dimensional data. A simple trick to overcome the associated curse of dimensionalty issue is to consider instead the distribution of intra-set distances (between pairs of real data points) and inter-set distances (between pairs of fake points). This metric was introduced in [yang:2020:evaluation_generative_models] for assessing the quality of generative models in symbolic music, but distances in that case are based on domain specific aspects of music such as rhythm or harmony, and cannot be easily extended to other types of data. A closely related metric based on maximum mean discrepancy [obray:2021:evaluation_metrics_graph], popular in the context of graph generative models, has been shown to be complicated due to its strong sensitivity to hyperparameters in the definition of the metric itself (leading to a sort of hyper-hyperparameters).
@@
10.2 Learning to (hyperparameter) optimize
In section 8.4 we covered connections between HPO and meta-learning from a problem-setting point of view, highlighting the connections between meta-parameters of a meta-learning algorithm and hyperparameters of a standard algorithm. In addition, one may also think of meta-learning an hyperparamter optimization algorithm, in that the task distribution is made of HPO problems, each described by a distribution over datasets and response functions of one or several standard learning algorithms. Samples from such a task distribution may involve various datasets and learning algorithms and traces of evaluations of hyperparameter configurations and their response value (e.g. validation error). The goal of the meta-learner in this context would be to learn a general strategy for efficiently searching the hyperparameter space based on data collected while optimizing hyperparameters across many different problems.
Early attempts use meta-data to warm start model-based methods, or learn loss curves; see [hutter2019automated, Chapter 2] for a survey. However, more recent works take this perspective one step further, motivated by successes of transformer-based language modelling and learning to optimize, a sub-field of meta-learning, [chen2022learning]. For example, [chen2022towards] proposed using a large language model to directly generate promising hyperparameter configurations based on a natural language description of the dataset and learning algorithm. The language model is fine-tuned on a corpus of hyperparameter optimization traces, allowing it to capture complex relationships between hyperparameters and adapt its search strategy to the problem at hand.
The usage of rich meta-data, past experience, and the capacity to leverage similarities across standard learning algorithms may hold the promise of discovering HPO algorithm that are more efficient than classic approaches. However, significant challenges remain in developing truly general and robust meta-learning for HPO, including handling the diversity of problem types, scalability to high-dimensional spaces, and transferability across vastly different domains and algorithms.
10.3 HPO and foundation models
@@
Zero-shot [wei2021finetuned] and in-context learning [brown2020language] have become standard paradigms for utilizing large language models (LLM), both of which require choosing context (or prompt) templates and in-context examples. LLMs’ performances at downstream tasks may be sensitive to such choices. These choices are often carried out manually by inspecting few rounds of interactions (prompt engineering). However, both problems can be formulated as hyperparamter optimization tasks, whereby the hyperparameters may be discrete choices over in-context examples or prompt templates, or even entire sequences over tokens. The resulting search space may be large and invariably discrete, and the response functions often involve several potentially expensive and potentially noisy calls to the LLM, which render the resulting HPO problems particularly challenging. Initial attempts include AutoPrompt [shin2020autoprompt], a gradient-based approach (cf. Chapter 7) that optimizes for discrete “trigger” tokens, Self-Adaptive In-Context Learning [wu2022self], which proposes a selection-then-rank framework for selecting instance-dependent examples; TRIPLE [shi2024efficient] which tackles template optimization under a best arm identification angle (algorithms proposed include successive halving; cf. Section 5.2.1); HbBoPs [schneider2024hyperband], which leverages a structure-aware Gaussian Process surrogate combined with Hyperband to efficiently select prompts in a black-box LLM setting; Black-box Discrete Prompt Learning [diao2022black] which uses a policy gradient estimator to optimize the prompt templates, among others. Further research is needed to study more efficient and scalable methods for optimizing prompts and in-context examples for large language models. Advances in reinforcement learning and gradient-based approaches (with discrete gradient estimators – see Section 7.6) may prove useful in these directions,
In addition, the emergence of agents and agent networks where various LLMs interact and act by accessing to external tools [schick2024toolformer, yang2024swe] provides another compelling application ground for HPO, in particular considering that many applications leverage closed-source LLMs queried through APIs. These agent-based systems introduce new hyperparameters to optimize, such as the selection and configuration of external tools available to the agents, including documentation, output formatting and error reporting, as well as the the communication protocols and the interactions graph between agents – see e.g. GPTSwarm [zhuge2024language]. Optimizing these hyperparameters presents unique challenges due to the complex interactions between agents, tools, and the underlying LLMs. Traditional HPO methods may need to be adapted or extended to handle the dynamic and potentially non-stationary nature of these systems. Furthermore, the use of closed-source LLMs via APIs introduces additional constraints, such as rate limits and usage costs, which must be considered in the optimization process.
Acknowledgements.
The work of PF was partially supported by Italian national project CAI4DSA, CUP B13C23005640006.